Thomas' Bike Shop stocks a high volume item that has a normally distributed demand during the reorder period. The average daily demand is 70 units, the lead time is 4 days, and the standard deviation of demand during the reorder period is 15.
a. How much safety stock provides a 95% service level to Thomas?
b. What should the reorder point be?
Answer #1

The average daily demand = 70 units

Lead time = 4 days

Mean lead time demand = The average daily demand * Lead time

= 70 * 4

= 280 units

The standard deviation of demand = 15 units

Ans a)

Desired service level = 95%

Z-value for achieving 95% service level = 1.65

Safety stock = Z * Standard deviation of demand

= 1.65 * 15

= 24.75 or 25 units

Ans b)

Appropriate re-order point = Mean lead time demand + Safety stock

= 280 + 25

= 305 units


answered by: Singge